(square root of x+11)=(square root of x)+1
Answer by ilana(307)

I'm not sure if this is right, but I'm guessing you meant:
(square root of (x+11))=(square root of x)+1.
So, I am going to write this sqrt(x+11)=(sqrtx)+1. Now, we want to isolate x on one side. Squaring both sides gives x+11=((sqrtx)+1)^2, so x+11=x+2sqrtx+1, so 2sqrtx=10, so sqrtx=5, so x=25. If we plug this value back into the problem, we do in fact get 6=6.